# On the Dimension of Pullback Attractors in Recurrent Neural Networks

### v1

@@ 2 Background
Roughly speaking two input sequences are close in this topology if they are close at only finite number of terms and maybe arbitrarily far away from each other at other terms. By Tychonoff’s theorem [21, Theorem 37.3], $\mathcal{U}$ is compact in this topology. Another topology on $\mathcal{U}$ is the uniform topology induced by the metric
@@
$C(\mathcal{U},X)$ is a complete metric space [21, Theorem 43.6].
@@ 2.1.1 Attractors, Entire Solutions and the Echo State Property
As explained in [22, Chapter 3], the behaviour of nonautonomous dynamical systems is best understood using an indexed family $\mathbf{A}=\{\mathbf{A}_{\mathbf{u}}\}_{\mathbf{u}\in\mathcal{U}}$ of subsets of $X$. For each $\mathbf{u}\in\mathcal{U}$, the set $\mathbf{A}_{\mathbf{u}}$ is called the $\mathbf{u}$-fiber of $\mathbf{A}$ and the set $\mathbf{A}$ is called a nonautonomous set. $\mathbf{A}$ is said to be non-empty (resp. compact) if each of the fibers are non-empty (resp. compact). Given a non-empty and compact nonautonomous set $\mathbf{A}=\{\mathbf{A}_{\mathbf{u}}\}_{\{\mathbf{u}\in\mathcal{U}\}}$, we say that $\mathbf{A}$ is $\Phi$-positively invariant if for any $\mathbf{u}\in\mathcal{U}$
@@
If a non-empty and compact nonautonomous set is $\Phi$-positively invariant, then there exist at least one entire solutions associated to every input sequence [22, Chapter 2, Lemma 2.20]. Note that, an entire solution, if it exist need not be unique. This generally is the case if for a given fixed $u\in U$, $g(x,u)$ has more than one fixed point. For example if for some $u\in U$ fixed, $z_{0},x_{0}\in X$ are fixed points of the map $g(x,u)$ with $x_{0}\neq z_{0}$, then there are infinitely many entire solution $\mathbf{x}$ one can assign to the input sequence $\mathbf{u}=\{u\}$ with $\mathbf{x}[0]=x_{0}$. In particular, while forward uniqueness is guaranteed for any well-define $g$, we may not always have backward uniqueness.
@@
is the Hausdorff semi-metric between two non-empty subsets $A$, $B$ of a metric space $(X,d)$. It can be easily checked that if $\text{dist}(A,B)=0$, then $\overline{A}\subseteq\overline{B}$ (See [28, Chapter 1]).
@@
It is important to remark that there is another definition of pullback and forward attraction suitable for one input sequence. These definitions are essentially equivalent to the pullback/forward attraction with respect to the process formulation. Since we are concern with all input sequences in $\mathcal{U}$, we use the definition of pullback and forward attraction suitable for this case (See [22, Chapter 3] for more details).
@@
$\mathbf{A}_{\mathbf{u}}=\{x\in X:\exists\>\text{an entire solution }\mathbf{x}\text{ associated to }\mathbf{u},n\in\mathbb{Z}\text{ satisfying }\>\mathbf{x}[n]=x\}.$
@@ 3.1 Pullback Attractor as the image of a continuous map
The proof will follow arguments of the proof of [15, Theorem 2.2.2]. Let $\mathbf{\Psi}:C(\mathcal{U},X)\to C(\mathcal{U},X)$ be the map $H\to\Phi(1,\theta^{0}(\cdot),H(\theta^{-1}(\cdot))$. We will show that $\mathbf{\Psi}$ is a $\mu$-contraction on $C(\mathcal{U},X)$. Let $F,H\in C(\mathcal{U},X)$. Then
@@ 3.2 Fractal Dimension of The Pullback Attractor
    if $M$ is a compact $m$-dimensional manifold, $\omega:M\to\mathbb{R}$ is a generic continuous function, $\phi:M\to M$ is a diffeomorphism and $\mathcal{V}$ is the set all of sequences of the form
@@
In the general case, $\text{dim}_{B}\>\mathcal{U}=\infty$ where $\mathcal{U}=U^{\mathbb{Z}}$ for some open subset $U$ of $\mathbb{R}^{N_{\text{in}}}$. As $\mathcal{H}_{\mathcal{U}}\subseteq\mathbb{R}^{N_{r}}$, we already know that $\text{dim}_{B}\>\mathcal{H}_{\mathcal{U}}\leq N_{r}$. In Theorem 3.8, we use the implicit functions theorem [1, Theorem 17.3] to show that $\mathcal{H}$ contains an open set and thus has dimension $N_{r}$. Hence however large one chooses $N_{r}$, the box-counting dimension of $\mathcal{H}$ will always be $N_{r}$. This result is perhaps not surprising since $H(\mathbf{u})$ depends on infinite past histories of $\mathbf{u}$ from the first term where $\mathbf{u}$ takes all possible values in an infinite dimensional metric space $\mathcal{U}=U^{\mathbb{Z}}$. Additionally, in proofs of universal approximations of echo state networks (ESNs) in the class of fading memory functionals, outline in [11], one tries to approximate to uniformly approximate input sequences by a finite dimensional projection using a neural network. The dimensionality of this projection is control by the number of output neurons of the neural network. The required functional on $\mathcal{U}$ is then approximated on the image of this finite dimensional projection. The dimensionality of this the image of this projection increases as the precision of the approximation increases.
@@
The following lemmas will be used in the proof. Lemma 3.9 is an immediate consequence of the submersion theorem [6, Corollary 2.4.9] and it states that submersions are open mappings. See also [29, Theorem 2.1] of the same result under weaker assumptions.
@@
The proof is an immediate consequence of [17, Theorem 1.2.17] which allows us to perturb the diagonal entries of $A$ to get a matrix whose determinant is non-zero and the result follows from the continuity of the determinant. ∎
@@ 3.2.1 Application to Echo State Networks
If $g$ is the state update equation of an ESN $i.e$ $g(u_{j},x)=\text{tanh}(Wx+W^{\text{in}}u_{j})$ where we assume that the connectivity matrices $W$ and $W^{\text{in}}$ both have full rank and $||W||_{2}<1$, then the assumptions of Theorem 3.8 holds. The connectivity matrices $W$ and $W^{\text{in}}$ are usually chosen randomly and thus the assumption of full rank holds almost surely. Here
$F(u,x)=\text{tanh}(W^{\text{in}}u_{2}+\text{tanh}(W^{\text{in}}u_{1}+\ldots(W^{\text{in}}u_{m}+Wx))\ldots)$
is differentiable since each of the partial derivatives exist and are continuous. In addition
@@
for all $(u_{j},x)\in U\times X$. The Jacobian with respect to $u$ is
$D_{u}F(u,x)$ $=\begin{bmatrix}\frac{\partial F_{1}(u,x)}{\partial u_{11}}&\frac{\partial F_{1}(u,x)}{\partial u_{12}}&\frac{\partial F_{1}(u,x)}{\partial u_{1N_{\text{in}}}}&\ldots&\frac{\partial F_{1}(u,x)}{\partial u_{21}}&\frac{\partial F_{1}(u,x)}{\partial u_{21}}&\ldots&\frac{\partial F_{1}(u,x)}{\partial u_{2N_{\text{in}}}}&\ldots&\frac{\partial F_{1}(u,x)}{\partial u_{mN_{\text{in}}}}\\ \vdots&\ddots&\ddots&\ldots&\ddots&\ddots&\dots&\ddots&\ldots&\vdots\\ \frac{\partial F_{N_{r}}(u,x)}{\partial u_{11}}&\frac{\partial F_{N_{r}}(u,x)}{\partial u_{12}}&\frac{\partial F_{N_{r}}(u,x)}{\partial u_{1N_{\text{in}}}}&\ldots&\frac{\partial F_{N_{r}}(u,x)}{\partial u_{21}}&\frac{\partial F_{N_{r}}(u,x)}{\partial u_{21}}&\frac{\partial F_{N_{r}}(u,x)}{\partial u_{2N_{\text{in}}}}&\ldots&\ldots&\frac{\partial F_{N_{r}}(u,x)}{\partial u_{mN_{\text{in}}}}\end{bmatrix}$
@@
$=\begin{bmatrix}S_{1}&S_{2}&&\ldots&S_{m}\end{bmatrix}\begin{bmatrix}W^{\text{in}}&\mathbf{0}&\ldots&\mathbf{0}\\ \mathbf{0}&W^{\text{in}}&\ldots&\mathbf{0}\\ \vdots&\ldots&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\ldots&W^{\text{in}}\end{bmatrix}$
where
@@
$S_{j}=\text{diag}(\operatorname{sech}^{2}(W^{\text{in}}u_{1}+Wy_{1}))\text{diag}(\operatorname{sech}^{2}(W^{\text{in}}u_{2}+Wy_{2}))\ldots\text{diag}(\operatorname{sech}^{2}(W^{\text{in}}u_{j}+Wy_{j}))$
and the $y_{j}$’s are defined in (17). Hence if we choose $m$ so that $mN_{\text{in}}\geq N_{r}$, we have that for all $u\in U^{m}$ and $x\in X$, $D_{u}F(u,x)$ is full rank. Thus by Theorem 3.8
$\text{dim}_{B}\mathcal{H}=N_{r}.$

### v2

@@ 2 Background
Roughly speaking two input sequences are close in this topology if they are close at only finite number of terms and maybe arbitrarily far away from each other at other terms. By Tychonoff’s theorem [23, Theorem 37.3], $\mathcal{U}$ is compact in this topology. Another topology on $\mathcal{U}$ is the uniform topology induced by the metric
@@
$C(\mathcal{U},X)$ is a complete metric space [23, Theorem 43.6].
@@ 2.1.1 Attractors, Entire Solutions and the Echo State Property
As explained in [24, Chapter 3], the behaviour of nonautonomous dynamical systems is best understood using an indexed family $\mathbf{A}=\{\mathbf{A}_{\mathbf{u}}\}_{\mathbf{u}\in\mathcal{U}}$ of subsets of $X$. For each $\mathbf{u}\in\mathcal{U}$, the set $\mathbf{A}_{\mathbf{u}}$ is called the $\mathbf{u}$-fiber of $\mathbf{A}$ and the set $\mathbf{A}$ is called a nonautonomous set. $\mathbf{A}$ is said to be non-empty (resp. compact) if each of the fibers are non-empty (resp. compact). Given a non-empty and compact nonautonomous set $\mathbf{A}=\{\mathbf{A}_{\mathbf{u}}\}_{\{\mathbf{u}\in\mathcal{U}\}}$, we say that $\mathbf{A}$ is $\Phi$-positively invariant if for any $\mathbf{u}\in\mathcal{U}$
@@
If a non-empty and compact nonautonomous set is $\Phi$-positively invariant, then there exist at least one entire solutions associated to every input sequence [24, Chapter 2, Lemma 2.20]. Note that, an entire solution, if it exist need not be unique. This generally is the case if for a given fixed $u\in U$, $g(x,u)$ has more than one fixed point. For example if for some $u\in U$ fixed, $z_{0},x_{0}\in X$ are fixed points of the map $g(x,u)$ with $x_{0}\neq z_{0}$, then there are infinitely many entire solution $\mathbf{x}$ one can assign to the input sequence $\mathbf{u}=\{u\}$ with $\mathbf{x}[0]=x_{0}$. In particular, while forward uniqueness is guaranteed for any well-define $g$, we may not always have backward uniqueness.
@@
is the Hausdorff semi-metric between two non-empty subsets $A$, $B$ of a metric space $(X,d)$. It can be easily checked that if $\text{dist}(A,B)=0$, then $\overline{A}\subseteq\overline{B}$ (See [31, Chapter 1]).
@@
It is important to remark that there is another definition of pullback and forward attraction suitable for one input sequence. These definitions are essentially equivalent to the pullback/forward attraction with respect to the process formulation. Since we are concern with all input sequences in $\mathcal{U}$, we use the definition of pullback and forward attraction suitable for this case (See [24, Chapter 3] for more details).
@@
$A_{\mathbf{u}}=\{x\in X:\exists\>\text{an entire solution }\mathbf{x}\text{ associated to }\mathbf{u},n\in\mathbb{Z}\text{ satisfying }\>\mathbf{x}[n]=x\}.$
@@ 3.1 Pullback Attractor as the image of a continuous map
The proof will follow arguments of the proof of [17, Theorem 2.2.2]. Let $\mathbf{\Psi}:C(\mathcal{U},X)\to C(\mathcal{U},X)$ be the map $H\to\Phi(1,\theta^{0}(\cdot),H(\theta^{-1}(\cdot))$. We will show that $\mathbf{\Psi}$ is a $\mu$-contraction on $C(\mathcal{U},X)$. Let $F,H\in C(\mathcal{U},X)$. Then
@@ 3.2 Fractal Dimension of The Pullback Attractor
Since assumption 3.3 amounts to uniform differentiability of $g$, it is sufficient for $g$ to be continuously differentiable and $X$ to be compact.
Remark 8.
@@
    if $M$ is a compact $m$-dimensional manifold, $\omega:M\to\mathbb{R}$ is a generic continuously differentiable function, $\phi:M\to M$ is a diffeomorphism and $\mathcal{V}$ is the set all of sequences of the form
@@
In the general case, $\text{dim}_{B}\>\mathcal{U}=\infty$ where $\mathcal{U}=U^{\mathbb{Z}}$ for some open subset $U$ of $\mathbb{R}^{N_{\text{in}}}$. As $\mathcal{H}_{\mathcal{U}}\subseteq\mathbb{R}^{N_{r}}$, we already know that $\text{dim}_{B}\>\mathcal{H}_{\mathcal{U}}\leq N_{r}$. In Theorem 3.8, we use the implicit functions theorem [1, Theorem 17.3] to show that $\mathcal{H}$ contains an open set and thus has dimension $N_{r}$. Hence however large one chooses $N_{r}$, the box-counting dimension of $\mathcal{H}$ will always be $N_{r}$. This result is perhaps not surprising since $H(\mathbf{u})$ depends on infinite past histories of $\mathbf{u}$ from the first term where $\mathbf{u}$ takes all possible values in an infinite dimensional metric space $\mathcal{U}=U^{\mathbb{Z}}$. Additionally, in proofs of universal approximations of echo state networks (ESNs) in the class of fading memory functionals, outline in [12], one tries to approximate uniformly approximate input sequences by a finite dimensional projection using a neural network. The dimensionality of this projection is control by the number of output neurons of the neural network. The required functional on $\mathcal{U}$ is then approximated on the image of this finite dimensional projection. The dimensionality of this the image of this projection increases as the precision of the approximation increases.
@@
The following lemmas will be used in the proof. Lemma 3.9 is an immediate consequence of the submersion theorem [6, Corollary 2.4.9] and it states that submersions are open mappings. See also [32, Theorem 2.1] of the same result under weaker assumptions.
@@
The proof is an immediate consequence of [19, Theorem 1.2.17] which allows us to perturb the diagonal entries of $A$ to get a matrix whose determinant is non-zero and the result follows from the continuity of the determinant. ∎
@@ 3.2.1 Application to Echo State Networks
If $g$ is the state update equation of an ESN $i.e$ $g(u,x)=\text{tanh}(Wx+W^{\text{in}}u)$ where $W\in\mathbb{R}^{N_{r}\times N_{r}}$ and $W^{\text{in}}\in\mathbb{R}^{N_{r}\times N_{\text{in}}}$. We assume that $u$ takes values in some bounded open set $U$. We want to check under which conditions the $m$-fold application map $F$ defined as
$F(u,x)=\text{tanh}(W^{\text{in}}u_{1}+W\text{tanh}(W^{\text{in}}u_{2}+W\text{tanh}(\ldots+W\text{tanh}(W^{\text{in}}u_{m}+Wx)))$
satisfies assumptions 3 and 2 of Theorem 3.8 for $m\in\mathbb{N}$ satisfying $m\geq N_{r}$. For assumption 3, it is sufficient that $||W||_{2}<1$. Indeed for any $(u_{j},x)\in U\times[-1,1]^{N}$,
@@
since $\text{sech}^{2}$ has an upper bound of $1$.
To investigate assumption 2, we assume that $F$ already satisfies assumption 3. Note that that this means that $F$ is a contraction and consider the input $v=(u,\ldots,u)\in U^{m}$ for some fixed $u\in U$. It follows that there exist a unique $z\in X$ satisfying $F(u,z)=z$. The Jacobian with respect to $u$ is
$D_{u}F(v,z)$ $=\begin{bmatrix}\frac{\partial F_{1}(v,z)}{\partial u_{11}}&\frac{\partial F_{1}(v,z)}{\partial u_{12}}&\frac{\partial F_{1}(v,z)}{\partial u_{1N_{\text{in}}}}&\ldots&\frac{\partial F_{1}(v,z)}{\partial u_{21}}&\frac{\partial F_{1}(v,z)}{\partial u_{21}}&\ldots&\frac{\partial F_{1}(v,z)}{\partial u_{2N_{\text{in}}}}&\ldots&\frac{\partial F_{1}(v,z)}{\partial u_{mN_{\text{in}}}}\\ \vdots&\ddots&\ddots&\ldots&\ddots&\ddots&\dots&\ddots&\ldots&\vdots\\ \frac{\partial F_{N_{r}}(v,z)}{\partial u_{11}}&\frac{\partial F_{N_{r}}(v,z)}{\partial u_{12}}&\frac{\partial F_{N_{r}}(v,z)}{\partial u_{1N_{\text{in}}}}&\ldots&\frac{\partial F_{N_{r}}(v,z)}{\partial u_{21}}&\frac{\partial F_{N_{r}}(v,z)}{\partial u_{21}}&\frac{\partial F_{N_{r}}(v,z)}{\partial u_{2N_{\text{in}}}}&\ldots&\ldots&\frac{\partial F_{N_{r}}(v,z)}{\partial u_{mN_{\text{in}}}}\end{bmatrix}$
@@
where $\frac{\partial F_{i}(v,z)}{\partial u_{jk}}$ is the partial derivative of $F_{i}$ with respect to the $k^{\text{th}}$ component of $u_{j}$, and
@@
$S_{j}=\begin{cases}\text{diag}(\operatorname{sech}^{2}(W^{\text{in}}u_{j}+Wy_{j}))&\quad\text{if}\>j=1\\ \prod_{i=1}^{j-1}\text{diag}(\operatorname{sech}^{2}(W^{\text{in}}u_{i}+Wy_{i})){W}&\quad\text{otherwise},\end{cases}$
and
$y_{i}=\text{tanh}(W^{\text{in}}u_{i+1}+Wy_{i+1})\quad\text{ for }i<m\>\text{ and }\>y_{m}=z.$
Taking
$D=\text{diag}(\operatorname{sech}^{2}(W^{\text{in}}v+Wz)),$
we have that
$D_{u}F(v,z)$ $=\begin{bmatrix}S_{1}W^{\text{in}}&|&S_{2}W^{\text{in}}&|&\ldots&|&S_{m}W^{\text{in}}\end{bmatrix}$
$=\begin{bmatrix}DW^{\text{in}}&|&DWW^{\text{in}}&|&\ldots&|&(DW)^{m-1}W^{\text{in}}\end{bmatrix}.$
To show that $D_{u}F(v,z)$ is full rank, it is sufficient to show that it has a $N_{r}\times N_{r}$ submatrix which is full rank. To this end, let $C\in\mathbb{R}^{N_{r}\times 1}$ be the first column of $W^{\text{in}}$ and consider the $N_{r}\times N_{r}$ matrix
$\begin{bmatrix}DC&|&DWC&|&\ldots&|&(DW)^{N_{r}-1}C\end{bmatrix}.$
Considering $W$, $W^{\text{in}}$ and $D$ as elements of $\mathbb{R}^{N_{r}^{2}}$, $\mathbb{R}^{N_{r}N_{\text{in}}}$ and $\mathbb{R}^{N_{r}}$ respectively, we prove the following stronger statement.
Theorem 3.12.
Given $W\in\mathbb{R}^{N_{r}\times N_{r}}$, $C\in\mathbb{R}^{N_{r}\times 1}$ and $D=\text{diag}(d_{1},\ldots,d_{N_{r}})\in\mathbb{R}^{N_{r}\times N_{r}}$, consider $(W,C,D)$ as an element of $\mathbb{R}^{N_{r}^{2}}\times R^{N_{r}}\times\mathbb{R}^{N_{r}}$. Then it is a generic property in $\mathbb{R}^{N_{r}^{2}}\times R^{N_{r}}\times\mathbb{R}^{N_{r}}$ that the matrix
$\begin{bmatrix}DC&|&DWC&|&\ldots&|&(DW)^{N_{r}-1}C\end{bmatrix}\in\mathbb{R}^{N_{r}\times N_{r}}$
is linearly independent.
Proof.
For a given $(W,C,D)\in\mathbb{R}^{N_{r}^{2}}\times R^{N_{r}}\times\mathbb{R}^{N_{r}}$, the determinant of the matrix
$\begin{bmatrix}DC&|&DWC&|&\ldots&|&(DW)^{N_{r}-1}C\end{bmatrix}$
is a polynomial in the entries $d_{i}$, $w_{ij}$ and $c_{j}$ of $D$, $W$ and $C$ respectively. If this polynomial does not vanish identically, its roots form a closed nowhere dense subset of $\mathbb{R}^{N_{r}^{2}}\times R^{N_{r}}\times\mathbb{R}^{N_{r}}$. We first proceed to show that there exist $(W,C,D)\in\mathbb{R}^{N_{r}^{2}}\times R^{N_{r}}\times\mathbb{R}^{N_{r}}$ for which the determinant of
$\begin{bmatrix}DC&|&DWC&|&\ldots&|&(DW)^{N_{r}-1}C\end{bmatrix}$
is non-zero. We take $D=\mathbb{I}_{N_{r}}$, $W=\text{diag}(w_{1},\ldots,w_{N_{r}})$ where the $w_{i}$’s are distinct and $C=(1,\ldots,1)^{T}$. Let $\alpha=(\alpha_{1},\ldots,\alpha_{N_{r}})\in\mathbb{R}^{N_{r}}$ and consider the linear combination
$\alpha_{1}DC+\sum_{j=2}^{N_{r}}\alpha_{j}(DW)^{j-1}C$ $=\alpha_{1}(1,\ldots,1)^{T}+\sum_{j=2}^{N_{r}}\alpha_{j}\sum_{k=1}^{N_{r}}w_{k}^{j-1}\mathbf{e}_{k}$
$=\sum_{j=1}^{N_{r}}\alpha_{j}\sum_{k=1}^{N_{r}}w_{k}^{j-1}\mathbf{e}_{k}$
$=\begin{bmatrix}1&w_{1}&w_{1}^{2}&\ldots&w_{1}^{N_{r}-1}\\ 1&w_{2}&w_{2}^{2}&\ldots&w_{2}^{N_{r}-1}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&w_{N_{r}}&w_{N_{r}}^{2}&\ldots&w_{N_{r}}^{N_{r}-1}\end{bmatrix}\begin{bmatrix}\alpha_{1}\\ \alpha_{2}\\ \vdots\\ \alpha_{N_{r}}\end{bmatrix}$
$=V\alpha$
where $\mathbf{e}_{k}$ is the $k^{\text{th}}$ canonical unit vector in $\mathbb{R}^{N_{r}}$. The matrix $V$ is a Vandermonde matrix [19, Chapter 0] whose determinant is given by
$\text{det}(V)=\prod_{1\leq i<j\leq N_{r}}(w_{i}-w_{j}).$
Since the $w_{i}$’s are distinct, we can conclude that $V$ is invertible. In particular $V\alpha=0$ if and only if $\alpha=0$. From this, we conclude that the columns of
$\begin{bmatrix}DC&|&DWC&|&\ldots&|&(DW)^{N_{r}-1}C\end{bmatrix}$
are linearly independent and thus the matrix has non-zero determinant. Thus the set of $W$, $D$ and $C$ for which
$\begin{bmatrix}DC&|&DWC&|&\ldots&|&(DW)^{N_{r}-1}C\end{bmatrix}$
is full rank form an open dense subset of $\mathbb{R}^{N_{r}^{2}}\times R^{N_{r}}\times\mathbb{R}^{N_{r}}$. ∎
Remark 9.
For $v=(u,\ldots,u)\in U^{m}$, we note that the fixed point $z$ depends continuously on $W$ and $W^{\text{in}}$. So if for a given $W$ and $W^{\text{in}}$, the Jacobian $D_{u}F(v,z)$ is not full rank, we can choose $(\widetilde{W},\widetilde{W}^{\text{in}},\widetilde{D})$ arbitrarily close to $(W,W^{\text{in}},D)$ so that the matrix
$\begin{bmatrix}DW^{\text{in}}&|&DWW^{\text{in}}&|&\ldots&|&(DW)^{m-1}W^{\text{in}}\end{bmatrix}$
is full rank. If $(\widetilde{W},\widetilde{W}^{\text{in}},\widetilde{D})$ is close enough to $(W,W^{\text{in}},D)$, we would have that $D_{u}\widetilde{F}(v,\widetilde{z})$ is also full rank where $\widetilde{F}$ redefines $F$ with the new matrices $\widetilde{W}$ and $\widetilde{W}^{\text{in}}$, and $\widetilde{z}$ is the new point in $\mathbb{R}^{N_{r}}$ satisfying $\widetilde{F}(v,\widetilde{z})=\widetilde{z}$. The reason for this is that in a sufficiently small neighborhood of a matrix, the rank of a matrix does not decrease.
4 Numerical Experiments
Here we test our theoretical result by driving echo state networks with $N_{r}=5,10,100$ hidden units with the trajectories of the Lorenz attractor [26] with the dynamics given by the ODE
$\dot{u}$ $=\sigma(v-u)$
$\dot{v}$ $=u(\rho-w)-v$
$\dot{w}$ $=uv-\beta w$
with $\rho=28,\sigma=10,\beta=8/3$ and the Rössler attractor with dynamics given by the ODE
$\dot{u}$ $=-v-w$
$\dot{v}$ $=u+av$
$\dot{w}$ $=b+w(u-c)$
with $a=b=0.1$ and $c=14$. We use numerical integration to generate the time $T$-map corresponding to $T=0.02$ and we sample $100$ initial conditions. All of the ESNs were allowed to run for $19000$ time steps and the state from time steps $19001$ to $20000$ recorded in a $100000\times N_{r}$ matrix. Notice that the long washout is to ensure that the trajectory is close to fibers of the corresponding pullback attractor (see remark 5).
[…]
Lorenz ($5$) $2.048$ $(2.018,2.077)$
Lorenz ($10$) $2.031$ $(2.001,2.062)$
Lorenz ($100$) $2.065$ $(2.049,2.08)$
Rössler ($5$) $1.771$ (1.735, 1.806)1.7351.806(1.735,\>1.806)( 1.735 , 1.806 )
Rössler ($10$) $1.8$ (1.78, 1.819)1.781.819(1.78,\>1.819)( 1.78 , 1.819 )
Rössler ($100$) $1.817$ (1.8, 1.834)1.81.834(1.8,\>1.834)( 1.8 , 1.834 )
[…]
Results of table 1 are remarkable. Although they confirm the upper bound in Theorem 3.5, they are strong evidence that the ESN embedding conjecture of Hart et al. of [17] is true in the general case. We leave investigating these further to a different paper.